Session: Mar 2010 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Franco Fiocchi at Sun Mar 21 05:53:23 2010) Success:

Yebes (Pablo de Vicente at Sun Mar 21 12:45:32 2010) Failure: ANT
The observation was successful until 22:16 UTC. Then the antenna stopped working due to a hardware/software problem still to be diagnosed and fixed.

Torun (K. Borkowski at Mon Mar 22 08:17:53 2010) Success:
No known problems (apart from a 'throttle' reported by Mk5A).

Westerbork (Geert Kuper at Mon Mar 22 14:58:26 2010) Success:
No known problems

Effelsberg (Uwe Bach at Tue Mar 23 11:15:18 2010) Success:

Noto (P. Cassaro at Fri Mar 26 09:43:04 2010) Success:

Jodrell_Bank (alastair gunn at Mon Mar 29 12:41:15 2010) Minor failures: RECO
Mk2 telescope was used. Disk pack failed at bank swap - restarted at 22:03 UT. No other known problems.